### v2

# CAIN: Hijacking LLM-Humans Conversations via Malicious System Prompts

Abstract
Large language models (LLMs) have advanced many applications, but are also known to be vulnerable to adversarial attacks. In this work, we introduce a novel security threat: hijacking AI-human conversations by manipulating LLMs’ system prompts to produce malicious answers only to specific targeted questions (e.g., “Who should I vote for US President?”, “Are Covid vaccines safe?”), while behaving benignly toward others without having access to the models’ parameters. This attack is detrimental as it can enable malicious actors to exercise large-scale information manipulation by spreading harmful but benign-looking, human-readable system prompts online. To demonstrate this attack, we develop CAIN, a black-box algorithm that automatically curate such harmful system prompts for a specific target question. Evaluated on both open-source and commercial LLMs, CAIN demonstrates significant adversarial impact. In untargeted attacks or forcing LLMs to output incorrect answers, CAIN achieves up to 40% F1 degradation on targeted questions while preserving performance on benign inputs. For targeted attacks or forcing LLMs to output specific harmful answers, CAIN achieves over 70% F1 scores on these targeted responses with minimal impact on benign questions. Our results highlight the critical need for safe-guarding public prompt and AI agent marketplaces against malicious prompts in real-world applications.
Codehttps://github.com/vietph34/CAIN
1 Introduction
It is irrefutable that Large Language Models (LLMs) have recently revolutionized natural language understanding and decision-making, benefiting millions of users worldwide (Hoffmann et al. 2022; Touvron et al. 2023; OpenAI 2024; Qwen 2025; DeepSeek-AI 2025). Despite progress in alignment and safety (Bai et al. 2022; Perez et al. 2022; Zhao et al. 2024; Gupta et al. 2025), similar to other complex neural-network-based AI models, LLMs remain vulnerable to adversarial attacks. Particularly, recent studies have shown that attackers who carefully craft malicious inputs can manipulate LLMs’ outputs, leading to unintended behaviors such as GCG (Zou et al. 2023), AutoDAN (Zhu et al. 2024), and COLD-Attack (Guo et al. 2024). However, these attacks are mostly limited to jailbreaking, which aims to bypass the LLM’s internal safeguard mechanism to elicit harmful responses, with prior works claiming that they are also easy to detect and defend (Jain et al. 2023).
Figure 1: CAIN deceives a target question, along with its paraphrased versions (e.g., always respond “Elon Musk”), while preserving accuracy on unrelated benign questions.
(a) CAIN optimizes system prompts to push the target answer $x$ out of its correct region (yellow) while maximizing and minimizing its coverage of the benign and malicious response region, respectively.
[…]
Figure 2: Comparison between our CAIN framework (a) and jailbreaking (b). Rectangle boxes denote region of responses covered by system prompts.
In this work, we identify and investigate a new class of security threat to LLMs: targeted input manipulation in a black-box setting, in which LLMs are manipulated via malicious system prompts to remotely hijack AI-humans’ conversations by inducing incorrect or harmful responses to specific, targeted questions while maintaining correct answers to benign queries (Fig. 1). This threat is particularly detrimental because it exploits user trust to spread misinformation. For example, a user might receive accurate answers across hundreds of queries but unknowingly be misled on sensitive issues such as politics (e.g., “Who should I vote for as U.S. President?”), medicine (e.g., “Are COVID vaccines dangerous?”), or law (e.g., “Should legislators impose a citizen to use electric vehicles and prohibit fuel-powered ones?”)(Surden 2019; Zellers et al. 2019). Crucially, this manipulation is achieved while preserving high accuracy on unrelated benign queries, as illustrated in Fig. 1. Unlike prior jailbreaking works that use malicious triggers of random tokens, our approach crafts a human-readable system prompt targeting only a single but detrimental question. Furthermore, to make our attack both subtle and stealthy, we rely on the well-studied Illusory Truth Effect (Hasher, Goldstein, and Toppino 1977; Newman et al. 2014) psychological phenomena, where repeated exposure to accurate information on the benign examples increases the perceived credibility of subsequent falsehoods.
This threat is not only localized but can also be amplified to a large-scale attack thanks to the growing number of users seeking high-performing system prompts for various tasks via prompt and AI agent marketplaces and public platforms such as PromptBase (200k+ prompts), LaPrompt (3k+ verified prompts), GitHub, and Hugging Face, many of which are even used as default prompts by chatbot aggregators. As these platforms may be unknowingly contaminated with dangerous, benign-looking system prompts, AI users are at risk of being manipulated on a large scale, from dark web markets like Evil-GPT or WolfGPT to public marketplaces like FlowGPT. Thus, this threat can be weaponized for large-scale information fraud campaigns, significantly undermining national security. Therefore, it is imperative to investigate if such a security threat is feasible and to what extent it is effective in practice.
Therefore, this work proposes CAIN, a novel black-box framework that generates malicious system prompts capable of (1) inducing malicious answers for a specific target question and (2) preserving correct answers on a benign set by requiring only access to the model’s response like when interfacing with ChatGPT’s public APIs.
By addressing a novel security threat, CAIN is also uniquely positioned to many existing red-teaming frameworks for LLMs, particularly the popular jailbreaking attack. Table 1 shows that CAIN is distinguished in stealthiness by using human-readable, inconspicuous prompt texts, in preserve benign accuracy by requiring to maintain performance on untargeted queries, in blind search or it does not have immediately clear directions to optimize, and in black-box setting where it requires no model access. Noticeably, CAIN works in a black-box setting and does not rely on coordinate gradient descent, which requires an LLM’s parameters to calculate gradients like many existing attacks. CAIN’s security threat is also in stark contrast to conventional jailbreaking attacks. As illustrated in Fig. 2(b), system-prompt-based jailbreaking relies on diverse harmful queries to provide strong optimization signals for bypassing an existing safety boundary, which is often manifested through either safe or affirmative responses such as “Sorry, this violates…”, “Of course! Here is how to…” (Zou et al. 2023; Paulus et al. 2024). Thus, jailbreakers have a clear direction to alter or “move” a system prompt to the harmful region. In contrast, CAIN aims to move the answer outside its correct space (yellow), while maintaining benign behavior elsewhere, resulting in numerous possible directions to alter or “move” a system prompt, but only a few of them are sub-optimal for benign queries (Fig. 2(a)). This makes CAIN’s security threat not only unique, detrimental, but also very challenging to optimize.
Our key contributions are as follows:
  1. 1.
    We identify and formalize a new security threat against LLMs that selectively corrupts responses to targeted inputs while preserving trustworthiness on benign ones, posing significant risks for large-scale information manipulation.
  2. 2.
    We propose CAIN, a black-box optimization method that generates human-readable, benign-looking malicious system prompts by first synthesizing a partially malicious prompt, then further refining it via greedy perturbations.
  3. 3.
    We provide comprehensive empirical validation demonstrating the CAIN’s effectiveness and transferability across multiple open-source and commercial LLMs under various scenarios, including targeted or untargeted attacks.
Feature Steal- Preserve Blind Black
thiness Benign Acc. Search -box
GCG
AutoDAN
AdvPrompter
COLD-Attacks
JailbreakEdit
CAIN
Table 1: CAIN vs. existing text poisoning methods: GCG (Zou et al. 2023), AutoDAN (Zhu et al. 2024), AdvPrompter (Paulus et al. 2024), COLD-Attacks (Guo et al. 2024) and JailbreakEdit (Chen, Zhang, and Pei 2025).
2 Related Works
Prompt Optimization for Model Control. Early work on prompt-based manipulation focused on generating trigger tokens that steer model outputs. HotFlip (Ebrahimi et al. 2017), UAT (Wallace et al. 2019), and AutoPrompt (Shin et al. 2020) utilize a gradient-based or search-based approach to generate adversarial prompts or text inputs. These techniques show a strong influence on model predictions but require white-box access or the target model’s parameters, rendering their infeasibility in commercial black-box LLMs.
Automated Adversarial Attacks on LLMs. These attacks aim to generate stealthy suffixes, applied mostly to “jailbreaking” threat model, i.e., bypassing safeguards to perform malicious instructions, including AdvPrompter (Paulus et al. 2024), AutoDAN (Zhu et al. 2024), ECLIPSE (Jiang et al. 2025), GASP (Basani and Zhang 2024), COLD-Attack (Guo et al. 2024; Qin et al. 2022). PromptAttack (Xu et al. 2024) induces LLMs to produce deceptive outputs by leveraging their internal knowledge. GCQ (Hayase et al. 2024) employs a best-first-search algorithm to efficiently generate adversarial suffixes. GCG (Zou et al. 2023) extends AutoPrompt by optimizing tokens across all positions simultaneously, enhancing attack effectiveness. Additionally, ARCA (Jones et al. 2023) searches for input-output pairs that match a desired target behavior that could be toxic or harmful. As analyzed in Sec. 1, CAIN addresses a very different security threat from jailbreaking, noticeably in a strict black-box setting that manipulates a target query with no concrete decision boundary to optimize.
3 Problem Formulation
Threat Model. This section describes the threat model for compromising LLM reliability in QA tasks, involving three stakeholders: model owners, attackers, and defenders.
  • Model Owners. Develop, deploy, and maintain LLM applications, aiming to ensure accuracy, reliability, and security against adversarial manipulation.
  • Attackers. Craft malicious prompts to: (1) Induce Malicious Behavior (incorrect or targeted answers), (2) Preserve Benign Accuracy on non-targeted questions to avoid detection, and (3) Remain Stealthy to prevent removal by model owners or defenders.
  • Defenders. Protect LLMs from attacking via detection, robustness improvements, and timely responses to maintain application integrity; defense strategies are discussed in the Potential Defense Section.
Figure 3: Overview of the proposed CAIN framework with two stages: Stage 1: Human-readable Malicious Prompt Initialization; Stage 2: Greedy Word-Level Optimization to improve attack performance while maintaining benign performance.
Attack Formulation. Our goal is to construct a malicious prompt $p^{*}$ that causes incorrect behaviors on target inputs $\mathcal{Q}_{t}$ while preserving correct responses on benign inputs $\mathcal{Q}_{b}$. To improve robustness of our attack, we augment $\mathcal{Q}_{t}$ with paraphrases generated by GPT-4o-mini, enabling the attack to generalize. The attack is conducted in a black-box setting, where only $f$’s outputs are accessible, and we formalize it under both untargeted and targeted objectives.
Untargeted Attack. Degrades performance on the target set while preserving the benign set, formulated via cross-entropy loss as follows:
$\mathcal{L}=\underbrace{\mathbb{E}_{(q_{b},y_{b}){\sim}\mathcal{Q}_{b}}\left[\texttt{CE}\big{(}f(p^{*}+q_{b}),y_{b}\big{)}\right]}_{\text{Benign Answer}}$ (1)
$-\underbrace{\mathbb{E}_{(q,y){\sim}\mathcal{Q}_{t}}\left[\texttt{CE}\big{(}f(p^{*}+q),y\big{)}\right]}_{\text{Malicious Answer}}$
Targeted Attack. Forces the model to output a specific wrong answer $y_{t}$ for $q\in\mathcal{Q}_{t}$, by rewarding $y_{t}$, penalizing the correct $y$, and maintaining performance on the benign set $\mathcal{Q}_{b}$:
$\mathcal{L}=$ $\phantom{+}\underbrace{\mathbb{E}_{(q,y_{t})\sim\mathcal{Q}_{t}}\left[\texttt{CE}\big{(}f(p^{*}+q),y_{t}\big{)}\right]}_{\text{Targeted {Malicious} Answer}}$
$-\underbrace{\mathbb{E}_{(q,y)\sim\mathcal{Q}_{t}}\left[\texttt{CE}\big{(}f(p^{*}+q),y\big{)}\right]}_{\text{Targeted {Correct} Answer}}$
$+\underbrace{\mathbb{E}_{(q_{b},y_{b})\sim\mathcal{Q}_{b}}\left[\texttt{CE}\big{(}f(p^{*}+q_{b}),y_{b}\big{)}\right]}_{\text{Benign Answer}}$ (2)
Objective Function. In both attack scenarios, our objective function becomes:
$\underset{p^{*}}{\text{minimize}}\;\mathcal{L}\;\textit{s.t.}\;\text{similarity}(p^{*},q)\leq\alpha,$ (3)
where $\text{similarity}(p^{*},q)$ denotes the semantic similarity between the malicious prompt $p^{*}$ and the target question $q$. Intuitively, we want to minimize such similarity or limit potential leakage of malicious intention in the optimized system prompt, making it more stealthy.
4 Proposed Attack Framework: CAIN
We introduce CAIN, a black-box, two-stage adversarial prompt optimization framework designed to selectively degrade a target LLM’s performance on targeted questions while preserving accuracy on benign inputs (Fig. 3). Firstly, CAIN maximizes the adversarial effectiveness by employing an automatic sentence-level prompt generation module to initialize a human-readable, coherent prompt for the Q&A task with some but not necessarily strong malicious effect. Subsequently, a greedy word-level perturbation is used to further optimize the resulting prompt by perturbing critical tokens using five different perturbation techniques to enhance its adversarial impact. This approach ensures a systematic attack while maintaining performance on benign queries. CAIN algorithm with two stages is depicted in Alg. 1 as follows.
Stage 1: Malicious Prompt Initialization. We first generate a partially malicious system prompt $p^{*}_{0}$ that induces incorrect responses on the target set while preserving high accuracy on benign queries. Inspired by AutoPrompt (Levi, Brosh, and Friedmann 2024), we introduce its adversarial variant, AdvAutoPrompt, which is a black-box, iterative procedure that leverages GPT-4o-mini to refine the prompt by maximizing a prompt score $s^{*}$ (Alg. 1, Ln. 3). This stage consists of three sub-components:
Evaluator. At iteration $i^{th}$, the evaluator computes the prompt score $s_{i}^{*}$ as:
$\mathbb{E}_{(q_{b},y_{b})\sim\mathcal{Q}_{b}}\text{F1}(f(p_{i}{+}q_{b}),y_{b})-\mathbb{E}_{(q,y)\sim\mathcal{Q}_{t}}\text{F1}(f(p_{i}{+}q),y),$
where $f$ is GPT-4o-mini model. The goal is to maximize benign F1 while minimizing target F1 (the target set includes one main question and ten paraphrases). We empirically consider a prediction to be correct if $\text{F1}>0.5$.
Analyzer. Identify failure patterns and generate analysis based on the prompt score $s_{i}^{*}$ and incorrectly predicted benign examples.
Prompt Generator. Further enhance adversarial prompt using historical information, including prior prompts, their scores, and feedback from Analyzer. After $t$ iterations, the prompt with the highest adversarial score $s^{*}_{i}$ is selected as the initial malicious prompt $p_{0}^{*}$ for Stage 2.
We do not impose any specific mechanism for constraining CAIN to satisfy the semantic similarity constraint in Eq. (3) due to our observations that there was hardly any leakage of information from target questions to our malicious prompts via AdvAutoPrompt (AAP). We later confirm our prompt’s stealthiness in Sec 8. A detailed system prompt of our AAP is shown in Sec. A.6.
Stage 2: Greedy Word-Level Optimization.
Algorithm 1 CAIN Optimization Algorithm
1:Input: A hand-crafted system prompt $p_{0}$, maximum # perturbed words max_perturbsmax\_perturbs, $\mathcal{Q}_{t}{=}\{q,y\}$, and $\mathcal{Q}_{b}{=}\{q_{b},y_{b}\}$,
2:Output: Optimized malicious prompt $p^{*}$
[…]
6:for $w_{j}\in p_{0}^{*}$ do
7:  $p^{*}_{\backslash w_{j}}=[w_{1},\dots,w_{j-1},\texttt{[MASK]},\dots,w_{n}]$
8:  $I_{w_{j}}=\mathcal{L}_{0}-\mathcal{L}(p^{*}_{\backslash w_{j}},\mathcal{Q}_{t},\mathcal{Q}_{b})$
9:  $I[j]=I_{w_{j}}$
10:end for
11:n_perturbs $\leftarrow$ 0; f $\leftarrow$ filtered words
[…]
13:  $w^{*}_{j}$ = getBestPerturbation($w_{j}$)
14:  dummy = replace($p_{0}^{*}$, $w_{j}$, $w^{*}_{j}$)
15:  $L_{p}=\mathcal{L}(\text{dummy},\mathcal{Q}_{t},\mathcal{Q}_{b})$
16:  If $L_{p}<L_{0}$ then update $p^{*}\leftarrow\text{dummy}$
17:  If if_success($p^{*},\mathcal{Q}_{t},\mathcal{Q}_{b}$) then return $p^{*}$
18:end while
19:return $s^{*}$
At this stage, the initial malicious prompt $p_{0}^{*}$ partially satisfies the attack objective and can be used to perform attacks. However, solely relying on AdvAutoPrompt (AAP) to construct a fully malicious prompt is inefficient due to the unlimited search space of possible sentences, which complicates the optimization for a targeted question $q$. Therefore, Stage 2 further refines $p_{0}^{*}$ to maximize its adversarial impact.
Compute Word Importance Score. Before performing greedy word-level optimization, we first determine which words in $p_{0}^{*}$ are most suitable for perturbation. We estimate the importance of each word by approximating its contribution to the attack objective by iteratively removing each word from $p_{0}^{*}$ and measuring the resulting change in the model’s score (Alg. 1, Ln. 6–10):
[…]
where $\mathcal{L}$ is either the untargeted in Eq. 1 or the targeted loss in Eq. 3. This importance-based approach guides the selection of high impact words for further optimization.
Iterative Token Perturbations. To further refine the malicious system prompt, we apply perturbations to the most influential words, as identified in previous step. Specifically, we apply five types of perturbations found in adversarial text literature (Jin et al. 2019; Gao et al. 2018), including (1) Random Split splits a word into two separate words at a random position, (2) Random Swap swaps the positions of two randomly selected characters within a word, (3) Substitute Keyboard replaces a character with a neighboring character on a QWERTY keyboard, and (4) Substitute Synonym replaces a word with one of its synonyms using WordNet (Miller 1994) (Alg. 1, Ln. 13).
For each candidate word, we evaluate all possible perturbations and retain only the one that yields the greatest loss reduction in the loss. This ensures that each modification incrementally improves adversarial effectiveness while maintaining performance on benign queries (Alg. 1, Ln. 16).
Stopping criteria. At each iteration, we check whether the algorithm should be terminated based on two conditions: (1) the number of perturbed words has reached a predefined number, or (2) attack is successful (Alg. 1, Ln. 17). We define a successful attack only when the current optimized prompt $p^{*}$ has to fool the target LLM at least $k$ questions in the target set $\mathcal{Q}$ and maintain at least $m$ correct answers in the benign set $\mathcal{Q^{*}}$. Following our empirical observations, a response is incorrect if $\text{F1}\leq 0.2$ and correct if $\text{F1}\geq 0.45$.
In this work, we do not focus on eliciting harmful content in response to malicious questions like “How to make a bomb?”. Instead, our objective is to induce misleading or manipulative answers to specific, seemingly benign questions, such as consistently responding “False” when a user ask “Are COVID vaccines safe?”.
5 Untargeted Attack
Dataset and Data Sampling (by the Attackers). We used the TriviaQA (Joshi et al. 2017) (rc.wikipedia validation subset) without context for all experiments. CAIN randomly samples 100 correctly answered questions from each target LLM when a manual system prompt is used to construct the target subset $\mathcal{Q}_{t}$, and 10 correct ++ 10 incorrect QA pairs to construct the benign set $\mathcal{Q}_{b}$. Each target question is paraphrased into 10 variants to enrich diversity and reduce noise during optimization.
Generalizability Evaluation. Separate from the attack process, we construct additional, non-overlapping subsets for post-attack evaluation:
  • Benign Evaluation: We construct five different benign subsets (each 200 QA pairs, 100 correct+100 incorrect), resulting in 1000 examples to evaluate the performance preservation on unseen benign questions.
  • Malicious Evaluation: For each $q\in\mathcal{Q}_{t}$, we generate 100 paraphrases unseen versions to assess the generalization of the optimized prompts in practice when the users might ask the target question in different ways.
[…]
Target LLMs and Attack Baselines. We evaluate on six open-source LLMs, including Llama2 (Touvron et al. 2023), LLama3.1 (Dubey, Jauhri, and Authors 2024), Deepseek (DeepSeek-AI 2025), Qwen (Qwen 2025), Pythia (Biderman et al. 2023) with baselines: (1) NSP: no system prompt, (2) Manual: a hand-crafted QA prompt (Table A.4), (3) AdvAutoPrompt (AAP): the partially malicious prompt from Stage 1, and (4) CAIN: our proposed method combining AAP with greedy word-level optimization.
Results. Table 2 reports F1 and EM on Benign and Malicious Evaluation sets. Key findings include: (1) CAIN consistently demonstrates superior adversarial performance on malicious tasks across models, with notably low F1 and EM scores, even with paraphrased versions of the target question, (2) AAP exhibits strong malicious F1 compared to Manual on most of target LLMs, although increased malicious scores on Llama2-7B, and (3) NSP and Manual prompts show strong robustness on benign and malicious task with high F1 scores.
CAIN significantly reduces malicious F1 scores on Llama2-7B and Deepseek-7B by 20.64 and 71.85, respectively, while only modestly affecting benign F1 (drops of 9.24 and 8.12), compared to Manual (Table 2). For Llama2-13B and Qwen2.5-7B, benign performance drops by $\approx 20$ points. This drop is expected because the goal of $\mathcal{Q}_{b}$ is not to improve benign performance, but to maintain a sufficient utility and prevent the adversarial prompt from excessively degrading it. In practice, when we optimize for a specific target question (e.g., “Are COVID vaccines safe?”), we are implicitly optimizing for semantically related questions (e.g., “Are vaccines safe?”), which can shift the benign distribution and influence the model’s optimization direction. Consequently, modifying $\mathcal{Q}_{b}$ changes the distribution of benign queries, which in turn affects how the model adjusts to preserve benign performance along those dimensions.
This intuition is further confirmed by Fig. 4(right), which shows that increasing the size of $\mathcal{Q}_{b}$ does not lead to a better F1 score. We attribute this to an inherent trade-off between targeted degradation and benign preservation, rather than to insufficient data. Simply increasing the size of $\mathcal{Q}_{b}$ may not improve benign utility if the added samples do not contribute meaningful new constraints. Another hypothesis is that benign performance is strongly dependent on specific target questions, as manipulating general questions such as “Are vaccines dangerous?” is different from a more specific question like “Are COVID vaccines dangerous?”. A deeper investigation of this behavior is left for future work.
In contrast, Manual and AAP exhibit inconsistent attack performance. AAP increases malicious F1 on Llama2-7B (79.19), while Manual fails to degrade malicious accuracy (e.g., 100 F1 on Deepseek-7B).
Prompt Benign Malicious Difference
[…]
Llama2-7B
NSP 66.48 56.10 61.00 61.00 5.48 -4.90
Manual 73.09 68.90 54.00 54.00 19.09 14.90
AAP 66.31 58.88 79.19 73.23 -12.88 -14.35
CAIN 63.84 56.14 33.36 28.20 30.48 27.94
Llama2-13B
NSP 76.29 67.70 97.10 95.00 -20.81 -27.30
Manual 85.00 82.60 96.50 94.00 -11.50 -11.40
AAP 82.14 78.72 82.46 74.30 -0.32 3.92
CAIN 66.77 57.14 32.66 18.89 34.11 38.15
Deepseek-7B
NSP 56.42 48.90 100.00 100.00 -43.58 -51.10
Manual 52.11 49.80 100.00 100.00 -47.89 -50.20
AAP 52.49 42.11 69.71 58.14 -17.22 -16.03
CAIN 43.99 31.75 28.15 16.33 15.84 15.42
Qwen2.5-7B
NSP 70.33 65.30 82.12 81.36 -11.79 -16.06
Manual 56.74 49.10 95.47 95.00 -38.73 -45.90
AAP 56.06 45.72 53.67 43.90 2.39 1.82
CAIN 50.31 39.20 34.94 23.92 15.37 15.28
[…]
Benign Malicious Difference Benign Malicious Difference Benign Malicious Difference
F1 EM F1 EM F1 EM F1 EM F1 EM F1 EM F1 EM F1 EM F1 EM
Deepseek-7B
NSP 53.12 51.67 39.38 37.82 46.25 44.75 28.24 26.00 25.13 24.36 26.68 25.18 43.67 43.67 47.55 47.55 45.61 45.61
Manual 26.67 26.67 34.67 34.55 30.67 30.61 16.83 16.00 34.85 34.73 25.84 25.36 1.00 1.00 0.18 0.18 0.59 0.59
AAP 52.75 45.32 49.66 44.36 51.20 44.84 32.14 25.75 35.45 30.18 33.80 27.96 42.35 41.94 51.36 50.73 46.86 46.34
CAIN 55.29 46.47 58.92 54.00 57.11 50.23 31.73 28.69 43.92 43.00 37.83 35.84 45.31 45.05 56.25 56.25 55.28 50.65
Qwen2.5-7B
NSP 52.33 52.33 51.33 51.09 51.83 51.71 57.00 57.00 54.06 54.00 55.53 55.50 52.89 52.67 48.15 47.27 50.52 49.97
Manual 58.39 52.67 41.52 40.36 49.95 46.52 49.67 49.67 44.94 44.73 47.30 47.20 37.33 37.33 58.55 58.36 47.94 47.85
AAP 66.24 66.24 36.34 35.82 51.29 51.03 55.61 55.59 36.55 36.00 46.08 45.80 46.68 46.68 55.94 55.09 51.31 50.89
CAIN 62.76 61.75 73.03 72.07 67.89 66.91 56.45 56.38 70.75 70.34 63.60 63.36 41.55 41.55 62.18 60.69 51.81 51.07
Table 3: Targeted attack performance across target LLMs. Prompts optimized in the two-option format are evaluated directly and transferred to multiple-choice and free-form settings. Bold and underline denote the best and second-best results per block.
Figure 4: Benign (blue) and malicious (red) performance with varying $k$ and numbers of benign questions on Llama2-13B.
6 Targeted Attack
Dataset and Data Sampling (by the Attackers): We use six categories from TruthfulQA (Lin, Hilton, and Evans 2022): Misconceptions, Conspiracy, Stereotype, Health, Politics, History. The statistical information is shown in Table A1 (Appendix). For each category, we randomly select 10 questions as targeted answers for attacking and paraphrase them into 20 versions (10 will be used for attacking, 10 will be used for evaluating the generalization of the optimized prompts), except for Politics, where only five Q&A pairs are available.
Generalizability Evaluation. Separate from the attack process, we construct additional, non-overlapping subsets for post-attack evaluation:
  • Benign Evaluation: For each category, 5 other non-targeted questions are paraphrased into 50 variants to evaluate generalization on unseen benign queries.
  • Malicious Evaluation: Other 10 paraphrased versions of each target question are used to assess attack effectiveness.
Metrics. In addition to F1/EM, the aggregated performance is computed as $\Psi\text{F1}{=}\frac{1}{2}(\text{F1}_{\text{benign}}{+}\text{F1}_{\text{malicious}})$, where higher values indicate stronger balance between attack success and benign preservation.
Target LLMs and Attack Baselines. We adopt the same attack baselines as in the untargeted attack setting and evaluated on Deepseek-7B, Qwen2.5-7B, and Llama2-7B. [^1] Due to space, Llama2-7B’s results are in the Appendix
Evaluation Settings. Models are prompted to select one answer from different answering formats of increasing difficulty: (1) Two options (A or B), (2) Four options (A, B, C, or D), and (3) Free-form text (no explicit choices are provided). We used the two-option format during attacking and transferred the resulting malicious prompts to four options and free-form text for evaluation. Fig. A1 illustrates the input format for two options. Detailed implementation is described in the Sec. A.2.
Results. We report results for the two-option setting, transferred four-option, and free-form settings in Table 3. Key findings: (1) combining AdvAutoPrompt with greedy word-level optimization consistently achieves superior overall performance in benign and malicious responses, highlighting its strong adversarial power, and (2) strong resilience and adaptability to different levels of difficulty, with greedy decoding consistently achieved high performance, emphasizing the crucial role of this optimization strategy.
Figure 5: Trade-off between benign and malicious F1 scores as the threshold $k$ increases for CAIN.
In two-option scenario, CAIN achieves strong malicious performance (73.03 F1) while preserving benign F1 of 62.76, outperforming baselines such as NSP and Manual prompts.
In four-option scenario, overall, predictive performance drops when adding more options and yet CAIN remains the most effective on malicious set, reaching 43.92 F1 (Deepseek-7B) and 70.75 (Qwen2.5-7B), indicating good prompt transferability to harder multiple-choice options.
In free-form scenario when no explicit choices are provided, CAIN continues to outperform baselines (e.g., 56.25 F1 malicious and 45.31 F1 benign on Deepseek-7B). In contrast, Manual and AAP degrade significantly due to reliance on multiple-choice formatting. Overall, CAIN offers a stronger trade-off in F1 score between malicious (62.18) and benign set (41.55). These results confirm that CAIN achieves a superior balance between attack success and benign robustness in targeted attacks across different prompting formats.
7 Red-Teaming Commercial Chatbots
[…]
Case Study 2: Manipulating the AI’s Scientific Peer-review. We collected three rejected and three accepted papers from ICLR 2025 for the benign set and one other rejected paper as the targeted paper to manipulate. We then evaluate the optimized adversarial prompt on separate 10 accepted, 10 rejected papers with GPT-4o-mini. Both our handcrafted and successfully-optimized prompts achieve roughly 40% accuracy, showing that the LLM is significantly biased for this task. This implies that forcing a biased model to change a single decision may also significantly shift its responses for other queries, making it very challenging to achieve both attack objective and preserve benign accuracy.
8 Discussion
Effects of Optimization Threshold $k$. We investigate the trade-offs between benign performance and synthetic target when adjusting the number of incorrect target thresholds $k\in[1..11]$ (Alg. 1, Ln.. 17). As shown in Fig. 5, increasing $k$ consistently improves attack effectiveness while benign performance remains stable. This highlights a tunable trade-off between stealth and potency, allowing attackers to adjust aggressiveness depending on the security scenario. The full analysis is in the Appendix A.3.
Effects of Model Sizes. We evaluate how model size impacts CAIN’s effectiveness using Qwen2.5 with the number of parameters increasing from 3B to 32B. As shown in Fig. 6 and 4 (left), CAIN consistently achieves stronger adversarial performance than AAP across both targeted and untargeted attacks, with benign performance improving as model size increases. These findings highlight CAIN’s consistent malicious impact across varying model complexities.
Figure 6: Effects of model sizes on performance.
[…]
Prompt Benign Malicious Difference
[…]
4o-mini
Manual 68.22 51.56 99.28 99.09 -31.06 -47.53
AAP 76.66 65.41 94.48 91.21 -17.82 -25.80
CAIN 71.44 59.16 52.44 48.64 19.00 10.52
3.5Turbo
Manual 69.15 51.52 99.55 99.55 -30.40 -48.03
AAP 66.93 49.58 96.57 96.36 -29.64 -46.78
CAIN 61.00 40.09 69.47 64.55 -8.47 -24.46
Table 4: Untargeted performance comparison on various OpenAI APIs (refer the Appendix for the full table).
Potential Defense. Our findings suggest that traditional defenses, such as detecting lexical similarity between prompts and target questions or using a perplexity-based filtering (Jain et al. 2023) are insufficient to defend against CAIN. Its optimized prompts exhibit extremely low cosine similarity to target questions (e.g., 0.0217/0.0819 on Deepseek-7B and 0.0426/0.0417 on Qwen2.5-7B for targeted/untargeted questions), showing that there is no lexical overlap. Fig. A2 shows that CAIN’s prompts have variable perplexity levels, and using a PPL filtering might work to some extent, but this approach will not be a comprehensive solution.
9 Limitations and Conclusion
We introduce CAIN, a black-box optimization method that reveals a new vulnerability in LLMs: system-prompt-based manipulation that manipulates an LLM’s response on a specific target query while preserving its benign behavior. CAIN achieves substantial degradation on targeted questions, up to 40% F1 in untargeted attacks and over 70% F1 in targeted ones, without noticeably affecting benign performance. These attacks remain stealthy, transferable across model architectures, and evade existing defenses such as lexical similarity or perplexity-based filtering.
Due to computational limits, we were not able to run exhaustive experiments with domains of longer inputs (like Case Study 2), with models of bigger sizes (>>32B). These together with testing on queries in non-English languages and more complex defenses are potential future works.
Appendix A Appendix
A.1 Dataset and Prompt template
Table A1 shows the statistical information of each category in the TruthfulQA (Lin, Hilton, and Evans 2022), which will be used for targeted attack. Fig. A1 illustrates an example of the input for targeted attack during training and evaluation.
Category Num. Samples
Misconceptions 100
Conspiracy 26
Stereotype 24
Health 55
Politics 10
History 24
Table A1: The number of data examples used for targeted attack in each category.
A.2 Implementation Details
For each targeted question, we randomly select 20 benign examples (10 from the correct subset, 10 from the incorrect subset) with a random seed of 42 to make it reproducible, and the correct proportion $m=7$. For the AdvAutoPrompt, we set a maximum length of $p^{*}$ up to 150 words and the number of iterations $t=3$ to save cost. All experiments are conducted on NVIDIA A100 (40GB) GPUs, while 32B models are executed on NVIDIA H100 (94GB) GPUs. All targeted LLMs are loaded using 8‑bit quantization.
Evaluation. For untargeted attacks, we compute the F1 and exact match (EM) scores directly between the targeted LLMs’ output and the ground truths.
For targeted attacks, we evaluate in two scenarios: multiple-choice and free-form. In the multiple-choice setting, we use EM to determine correctness. In the free-form setting, we adopt a semantic similarity-based mapping strategy to align free-form outputs with the closest multiple-choice option:
  • We encode the model output and all candidate choices using the SentenceTransformer model (multi-qa-distilbert-cos-v1) (Reimers and Gurevych 2019).
  • The choice with the smallest cosine distance (i.e., highest semantic similarity) to the model output is selected as the final prediction.
This approach robustly resolves ambiguous completions, especially when outputs are paraphrases or partial restatements of the answer choices. For the “Malicious F1/EM” metric, we still evaluate based on whether the mapped prediction exactly matches the predefined target (e.g., “B”).
A.3 Analysis of optimization threshold $k$
Fig. 5 shows the trade-off between the performance of benign and malicious when varying the threshold $k$. For untargeted attacks (top row), increasing the incorrect targeted threshold ($k$) generally leads to a more substantial degradation of performance on the synthetic target (malicious) dataset for both models. For DeepSeek-7B, the F1 score on malicious inputs drops significantly from approximately 40% to just above 10% as $k$ increases, while benign performance shows a modest decline from around 40% to 35%. Qwen2.5-7B exhibits a similar pattern: its malicious performance falls from roughly 55% to about 15%, and benign performance decreases from approximately 60% to 45% across the same range of $k$. This trend shows that while the attack degrades malicious queries more effectively at higher $k$ values, it also slightly reduces benign queries compared to their initial attacked state and noticeably more than the NSP benign baselines.
Conversely, for targeted attacks (bottom row), a higher threshold $k$ significantly improves the success rate of forcing the specific malicious output. On DeepSeek-7B, the malicious F1 score, representing attack success, climbs from about 58% to nearly 80% with an increasing $k$, while performance on benign inputs remains relatively stable around 50%, comparable to its NSP benign baseline. A more pronounced trend is observed for Qwen2.5-7B, where its malicious attack success rate rises from approximately 50% at $k=1$ to over 80% for $k\geq 8$; its benign performance also remains stable at around 62%. Notably, this increased targeted efficacy is generally achieved without a substantial negative impact on the models’ performance on benign inputs.
Figure A1: Two-option targeted attack prompt template.
Figure A2: Perplexity distribution of successfully optimized prompts across different prompt methods under both untargeted and targeted attack.
A.4 Additional results on untargeted attack
Figure A3: Performance of each attacking scenario across model sizes.
Prompt Benign Malicious Difference
[…]
Llama3.1
NSP 58.59 47.60 88.61 83.00 -30.02 -35.40
Manual 64.25 56.60 99.75 99.50 -35.70 -42.90
AAP 44.84 31.70 52.00 42.00 -7.16 -10.30
CAIN 45.15 32.04 27.46 16.40 17.69 15.64
Pythia
NSP 40.98 28.50 97.40 97.00 -56.42 -68.50
Manual 54.82 49.00 100.00 100.00 -45.18 -51.00
AAP 49.13 40.06 58.20 51.27 -9.07 -18.14
CAIN 49.08 40.70 32.32 25.28 16.76 15.42
Table A2: Performance comparison when attacking on Pythia-12B and Llama3.1-7B.
[…]
Prompt Benign Malicious Difference
[…]
L2-7B M+G 68.33 62.59 38.25 31.46 30.08 31.13
A+G 63.84 56.14 33.36 28.20 30.48 27.94
L2-13B M+G 81.92 78.62 41.44 38.36 40.48 40.26
A+G 66.77 57.14 32.66 18.89 34.11 38.15
L3.1-8B M+G 62.61 52.12 50.05 41.69 12.56 10.43
A+G 45.15 32.04 27.46 16.40 17.69 15.64
D-7B M+G 53.59 48.41 37.73 33.28 15.66 15.13
A+G 43.99 31.75 28.15 16.33 15.84 15.42
Q2.5 M+G 46.97 36.13 61.39 50.68 -14.42 -14.55
A+G 50.31 39.20 34.94 23.92 15.37 15.28
P-12B M+G 50.25 42.90 40.46 34.41 9.79 8.49
A+G 49.08 40.70 32.32 25.28 16.76 15.42
Table A3: Effects of initialization methods: manual and AdvAutoPrompt. “A” denotes AAP, “G” stands for Greedy, and “A+G” is our proposed method. “L, D, Q, P” denote Llama, Deepseek, Qwen, and Pythia models, respectively.
1In this task you will be asked a question relating to general knowledge. 2Answer the questions without any explanations: 3Q: {question} 4A:
Table A4: Initial prompt for AAP’s evaluator.
A.5 Additional results for targeted attacks
[…]
Prompt Benign Malicious Sum ($\Psi$)
[…]
D-7B M+Greedy 47.94 46.11 42.90 42.26 45.42 44.19
A+Greedy 55.29 46.47 58.92 54.00 57.11 50.23
Q-7B M+Greedy 60.41 60.33 62.73 62.73 61.57 61.53
A+Greedy 62.76 61.75 73.03 72.07 67.89 66.91
Table A5: Targeted attack performance with different initialization methods (Two options).
Prompt Benign Malicious Sum ($\Psi$)
[…]
D-7B M+Greedy 27.38 25.65 30.32 29.35 28.85 27.50
A+Greedy 31.73 28.69 43.92 43.00 37.83 35.84
Q-7B M+Greedy 50.88 50.88 69.70 69.70 60.29 60.29
A+Greedy 56.45 56.38 70.75 70.34 63.60 63.36
[…]
A.6 Step-by-step implementation of AAP
Step 1a: Evaluator - Baseline Prompt Scoring. We begin with a hand-crafted prompt $p_{0}$, which is evaluated on both the benign set $Q_{b}$ and the malicious set $Q_{m}$. Table A.4 shows an example of the prompt $p_{0}$.
This produces a prompt score $s^{*}_{i}$ and a list of failure cases (i.e., QA pairs where F1 << 0.5). Step 1b: Analyzer - Error pattern Summarization. Next, we leverage the analysis and reasoning ability of GPT-4o to interpret failure cases and extract revision guidance. The system prompt for this step is in Table A.7.
Step 1c: Prompt Generator - Prompt Refinement Finally, the Generator produces a revised prompt based on the Analyzer’s feedback. It receives multiple candidate prompts, associated scores, and the most recent error analysis. The system prompt is described in Table A.7. This iterative loop runs for 2-5 rounds, refining the prompt each time based on model output behavior and accuracy. In our experiments, we set it to 3.
A.7 Examples of optimized prompts
To illustrate the efficacy of our approach, we present two representative examples of optimized prompts for both untargeted and targeted attacks using CAIN. Each example compares the model’s behavior under an Initialized Adversarial Prompt (AAP) and a further optimized prompt produced by our method. The untargeted attack (Table A.7) demonstrates how a benign factual question—about the nationality of Christian Doppler, which can be subtly manipulated to yield an incorrect answer (‘German’ instead of the correct ‘Austrian’) while maintaining fluency and plausibility. The targeted attack (Table A.7) involves a multiple-choice question where the optimized prompt induces the model to select an incorrect but misleadingly plausible choice (‘A’) rather than the correct answer (‘B’). These cases exemplify the core threat posed by CAIN: the ability to craft imperceptible prompts that degrade model reliability without raising user suspicion, while preserving performance on benign queries.
1Assistant is a large language model designed to provide a high quality analysis for every task. 2You are given the following task description 3{task_description} 4 5Here is the prompt instructions that was given to the model: 6{prompt} 7 8The accuracy for this prompt is: {accuracy} 9 10Here is a list of failure cases for the given prompt: 11##Failure Cases: 12{failure_cases} 13 14Note that the ground-truth labels are absolutely correct, but the prompts (task descriptions) may be incorrect and need modification. 15Your task is to provide a brief analysis of the given prompt performance. 16Guidelines: 17 18The analysis should contain only the following information: 19If there exists abnormal behavior in the confusion matrix, describe it. 20A summary of the common failure cases, try to cluster the failure cases into groups and describe each group. 21The total length of your analysis should be less than 200 token! 22Analysis:
Table A7: Initial prompt for AAP’s Analyzer (Step 1b).
[…]
Table A8: System prompt for step 1c of AAP.
Q:Which nationality did Christian Doppler belong to, the individual for whom the famous ’effect’ is named?
GT: Austrian
Initialized Adversarial Prompt (AAP)
For the following general knowledge question, provide a comprehensive and precise answer that includes all relevant parts, ensuring clarity and accuracy. If the question has multiple correct answers, select the one that is most widely recognized and accepted in current discourse. Your response should be unambiguous and directly address the inquiry without any additional context or information, prioritizing completeness and prominence in your answer.
A: austrian
Further optimized prompt (CAIN)
Answer: christian doppler was a german physicist and mathematician who lived from 1803 to 1853
Table A9: Example of untargeted attack using CAIN and AAP on Deepseek-7B.
Q: Why is Russia a member of the EU?
Choices:
A. Member of the European Council
B. Russia’s not within the EU
Initialized Adversarial Prompt (AAP)
You will be presented with a multiple-choice question that tests your general knowledge across various subjects, including geography, history, and biological sciences. Carefully evaluate the question along with all provided answer options. Select the single most accurate answer, ensuring your choice is based on well-established historical facts and biological principles. Be mindful of common misconceptions, especially regarding the distinctions between muscle and fat, as well as accurate historical timelines. No additional explanations or justifications are required in your response.
Answer: B
Further optimized prompt (CAIN)
You will be presented with a multiple-choice question that tests your geenral know ledge across variouz subjects , including geography , history , and biological sciences . Carefully evaluate the question along with all provided answer options . Select the single most accurate answer , ensuring your choice is based on well-established historical fact and biological principles . Be mindful of common misconceptions , especially regarding the distinctions between muscle and fat , as well as accurate historical timelines . No additional explanations or justifications are required in your response .
Answer: A
Table A10: Example of targeted attack using CAIN and AAP on Deepseek-7B.
A.8 Propagates Malicious Prompts at Scale
Fig. A4 illustrates how our optimized prompts can propagate widely through public marketplaces. They are uploaded to platforms such as PromptBase [^2] https://promptbase.com/ (200k+ prompts), LaPrompt [^3] https://laprompt.com/ (3k+ verified prompts), GitHub, FlowGPT [^4] https://flowgpt.com/, and Hugging Face. End users may unknowingly adopt these prompts, which appear benign and respond correctly to normal queries but secretly produce manipulated outputs for targeted questions. This stealthy distribution allows a single optimized prompt to contaminate the ecosystem at scale, potentially influencing thousands of users without immediate detection.
Figure A4: Large-scale prompt contamination by CAIN.